OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No Solution

OpenStudy (anonymous):

@gitahimart

OpenStudy (anonymous):

first distribute the -3 into the absolute term: \[\left| -6x-18 \right| = -12\] so this means that either: -6x-18 = -12 or 6x+18 = -12 can you solve for the x-values now?

OpenStudy (anonymous):

I Got C Is That The Answer?

OpenStudy (anonymous):

-6x = -12+18 -6x =6 x =-1 or 6x+18 = -12 6x =-12-18 6x=-30 x=-5 so x values should be -1 and -5, so the second choice is correct

OpenStudy (anonymous):

-6x-18=-12 -6x=-12+18 x=-6/6 =-1

OpenStudy (anonymous):

therefore x=-1

OpenStudy (anonymous):

@Deldrickg did u understand

OpenStudy (anonymous):

Yes.